# Gridlines Mitigate Sine Illusion in Line Charts

### v1

@@ Abstract
Sine illusion happens when the more quickly changing pairs of lines lead to bigger underestimates of the delta between them. We evaluate three visual manipulations on mitigating sine illusions: dotted lines, aligned gridlines, and offset gridlines via a user study. We asked participants to compare the deltas between two lines at two time points and found aligned gridlines to be the most effective in mitigating sine illusions. Using data from the user study, we produced a model that predicts the impact of the sine illusion in line charts by accounting for the ratio of the vertical distance between the two points of comparison. When the ratio is less than 50%, participants begin to be influenced by the sine illusion. This effect can be significantly exacerbated when the difference between the two deltas falls under 30%. We compared two explanations for the sine illusion based on our data: either participants were mistakenly using the perpendicular distance between the two lines to make their comparison (the perpendicular explanation), or they incorrectly relied on the length of the line segment perpendicular to the angle bisector of the bottom and top lines (the equal triangle explanation). We found the equal triangle explanation to be the more predictive model explaining participant behaviors.
@@ Introduction
First formally introduced by Cleveland and McGill in 1984 [7], the sine illusion describes a perceptual error involving a pair of lines with slope changing at an increasing rate. The more quickly changing pairs of lines can lead to bigger underestimates of the delta between them. This illusion is common in the real world. For example, multi-line line charts, stream graphs, and area charts all harbor opportunities for this bias, as viewers are often required to compare vertical distances to draw conclusions [28].
When comparing the length of line segments, the sine illusion happens as viewers rely on non-relevant visual cues, which can take several forms. For example, Day et al. [9] posit that the sine illusion can be explained as a perceptual comprise of the greater overall distance between the two lines around the points of comparison, similar to drawing out a circle around each point of comparison. Viewers consider the general area surrounding the points of comparison similar to viewing a Müller-Lyer line illusion [9, 15] and the ‘hull area’ proxy presented by Jardine et al. [16]. Others suggest that, instead of using an area proxy, viewers could be relying on the orthogonal distance (or the minimal distance) between the two lines instead of the vertical distance to make the delta comparisons [28, 4], similar to doing a Deming regression [23]. In this paper, we test these two explanations and create two models to determine which one might be a closer explanation to sine illusions. This allows us to infer what people were actually considering during the line segment comparison task. We refer to the two models as the ‘perpendicular model’ and the ‘equal triangle’ model, which we describe more closely in Section 2.7.
@@ 2.8 When Time 1 is Actually Larger
We also examine how the three design conditions influenced response accuracy when the vertical distance between the two lines at Time 1 is larger than that at Time 2. In this case, relying on the sine illusion would help the participants. We found the opposite effect, such that participants performed worse with the gridlines condition and best with the dotted lines condition. When they viewed the dotted line charts, their accuracy was 4.18 times higher than the gridlines condition (p < $0.001$). When they viewed the default normal condition, their accuracy was 2.75 times higher than the gridlines condition (p < $0.001$). However, despite this reversal in condition effectiveness, participants overall performed significantly better (t = $61.84$, p < $0.001$) in the scenario where Time 1 delta is actually larger (ratio of Time 1 over Time 2 is greater 1, $M_{accuracy}$ = $0.89$), compared to the scenario where Time 1 delta is smaller (ratio of Time 1 over Time 2 is smaller than 1, $M_{a}ccuracy$ = $0.61$). See annotation on the left-most panel in Figure 3.
We turn to the perpendicular and equal triangles models to better understand what might have driven this result. While further investigation is needed to draw causal conclusions, we suspect that the existence of the gridlines made participants double guess their response instead of relying on the sine illusion, which decreased accuracy. When the gridlines are present, participants can compare the length of the gridlines between the two lines at Time 1 and Time 2 to make their decisions. But length comparison is subjected to imprecision and perceptual bias, as length encoding is not an extremely precise encoding channel [18, 5, 6]. As a result, they make mistakes and perform with lower accuracy. However, when the gridlines are turned off, participants tend to rely on the sine illusion and compare the equal triangle distance for the task, which tends to result as Time 1 being perceived as larger. Thus when Time 1 delta is actually larger, relying on the sine illusion, rather than making a length comparison, results in a higher overall accuracy across all conditions, and a relatively lower accuracy for gridlines.
@@ 3 Design Guidelines and Summary of Findings
We produced a model that predicts the likelihood and severity of the sine illusion in line charts based on the ratio of the vertical distance between the two points of comparison. We found that, in general, accuracy drops below chance when the ratio falls above 0.7 for default line charts, and below 75% when the ratio falls above 0.5. This means that when participants are comparing the difference between two lines at two different time points, if that difference is less than 50%, then people will struggle. And if that difference is less than 30%, people will struggle significantly.
@@
In general, we found that people make mistakes in this task because they are not comparing the vertical distance between the two lines. They are actually comparing the length of the line segments perpendicular to the angle bisector of the bottom and top lines, as modeled by the equal triangle heuristics.

### v2

@@ Abstract
The sine illusion is an underestimation of the difference between two lines when both lines have increasing slopes. We evaluate three visual manipulations on mitigating sine illusions: dotted lines, aligned gridlines, and offset gridlines via a user study. We asked participants to compare the deltas between two lines at two time points and found aligned gridlines to be the most effective in mitigating sine illusions. Using data from the user study, we produced a model that predicts the impact of the sine illusion in line charts by accounting for the ratio of the vertical distance between the two points of comparison. When the ratio is less than 50%, participants begin to be influenced by the sine illusion. This effect can be significantly exacerbated when the difference between the two deltas falls under 30%. We compared two explanations for the sine illusion based on our data: either participants were mistakenly using the perpendicular distance between the two lines to make their comparison (the perpendicular explanation), or they incorrectly relied on the length of the line segment perpendicular to the angle bisector of the bottom and top lines (the equal triangle explanation). We found the equal triangle explanation to be the more predictive model explaining participant behaviors.
@@ Introduction
First formally introduced by Cleveland and McGill in 1984 [7], the sine illusion describes a perceptual error where more quickly changing pairs of lines can lead to bigger underestimates of the delta between them. Researchers printed out vertical lines of equal length following a sine wave pattern [9] and noticed that the lines at the straight oblique sections appeared shorter than those in the turns. This illusion is common in the real world. Multi-line line charts, stream graphs, and area charts all harbor opportunities for this bias, whenever viewers compare vertical distances between lines [28].
When comparing the deltas, the sine illusion happens as viewers rely on one of many potentially irrelevant visual cues as proxies for the actual distance between two lines For example, viewers might compare the areas of the regions near the two points of comparisons between the two lines rather than just the vertical distances, similar to the Müller-Lyer line illusion [9, 15], where viewers include the arrow tips in the length comparison to overestimate the line length, and the ‘hull area’ proxy presented by Jardine et al. [16], where the viewer perceives an implied hull bounded by the bars when comparing the means of bar sets. Others suggest that, instead of relying on the vertical distance between two points with a common X-axis value, viewers might rely on the orthogonal distance (or the minimal distance) which lead to non-veritical comparisons between points on the two lines that do not share a common x-axis value [28, 4], similar to doing a Deming regression [23]. We contrast these two explanations by pitting two models in competition to best model the sine illusion. We refer to the two models as the ‘perpendicular model’ and the ‘equal triangle’ model, which we describe more closely in Section 2.7.
@@ 2.8 When Time 1 is Actually Larger
When the vertical distance between the two lines at Time 1 is larger than that at Time 2, relying on the sine illusion would help the participants. In these cases, participants performed worse with the gridlines condition and best with the dotted lines condition. When they viewed the dotted line charts, their accuracy was 4.18 times higher than the gridlines condition (p < $0.001$). When they viewed the default normal condition, their accuracy was 2.75 times higher than the gridlines condition (p < $0.001$). However, despite this reversal in condition effectiveness, participants overall performed significantly better (t = $61.84$, p < $0.001$) in the scenario where Time 1 delta is actually larger (ratio of Time 1 over Time 2 is greater 1, $M_{accuracy}$ = $0.89$), compared to the scenario where Time 1 delta is smaller (ratio of Time 1 over Time 2 is smaller than 1, $M_{a}ccuracy$ = $0.61$). See annotation on the left-most panel in Figure 3.
We turn to the perpendicular and equal triangle models to better understand what might have driven this result. While further investigation is needed to draw causal conclusions, we suspect that the existence of the gridlines made participants double-guess their response instead of relying on the sine illusion, which decreased accuracy. When the gridlines are present, participants can compare the length of the gridlines between the two lines at Time 1 and Time 2 to make their decisions. But length comparison is subjected to imprecision and perceptual bias, as length encoding is not an extremely precise encoding channel [18, 5, 6]. As a result, they make mistakes and perform with lower accuracy. However, when the gridlines are turned off, participants tend to rely on the sine illusion and compare the equal triangle distance for the task, which tends to result as Time 1 being perceived as larger. Thus when Time 1 delta is actually larger, relying on the sine illusion, rather than making a length comparison, results in a higher overall accuracy across all conditions, and a relatively lower accuracy for gridlines.
@@ 3 Design Guidelines and Summary of Findings
We produced a model that predicts the likelihood and severity of the sine illusion in line charts based on the ratio of the deltas between the two points of comparison. We found that, in general, accuracy drops below chance when the ratio falls above 0.7 for default line charts, and below 75% when the ratio falls above 0.5. This means that when participants are comparing the difference between two lines at two different time points, if that difference is less than 50%, then people will struggle. And if that difference is less than 30%, people will struggle significantly.
@@
In general, people make mistakes in this task because they are not comparing the vertical distance between the two lines. They are comparing the length of the line segments perpendicular to the angle bisector of the bottom and top lines, as modeled by the equal triangle heuristics.